A tree 33 m tall and 0.55 m in diameter at its base produces about 67000 L of oxygen per year. A person who exercises vigorously several hours a day might take in about 22 m of all per day. How many trees are needed to provide oxygen to 900 people breathing?
1
Expert's answer
2019-09-12T03:33:33-0400

Person consumes "22m^3" air per day="0.2\times 22=4.4m^3\ O_2\\"

"\because\ air\ contains\ 20\ percentage \ of \ oxygen"

900 people will need "4.4\times 900=3960m^3\ O_2" per day

total oxygen required per year = "3960\times 365=1445400m^3\ O_2"

Amount of Oxygen given by 1 tree per year = "67000\ l=67\ m^3"

Total no. of trees required = "\frac{1445400}{67}=21573.134"

So, nearly 21573 trees will be required.